# First-principles design for strain-tunable exciton dynamics in 2D materials

### v1

@@ Results
To conclude, we have developed a fully first-principles-based semiclassical framework for modeling the time evolution of macroscopic distributions of quasiparticle distributions on inhomogeneously strained monolayer semiconductors. Applied to strained WS2 monolayers with idealized polynomial strain profiles, our model captures the qualitative impact of band curvature and energy gradients on exciton drift and diffusion, in the absence of scattering and relaxation mechanisms. The model qualitatively reproduces experimentally observed behaviors such as drift toward strain maxima, anisotropic broadening, and negative diffusion. It also provides a physical basis for estimating the ballistic mean free time in idealized regimes. These results offer fundamental insight into how strain modulates exciton transport in 2D semiconductors and establish a foundation for ab initio-informed design principles in straintronic devices, where strain landscapes are engineered to control quasiparticle motion. The model allows for straightforward incorporation of exciton interactions, enabling the method to evolve toward predictive modeling of realistic transport in complex strain landscapes.
@@
Acknowledgments: We thank Moshe Haratz, Lev Melnikovski, and Galit Cohen for valuable discussions. Computational resources were provided by the ChemFarm local cluster at the Weizmann Institute of Science and the Max Planck Computing and Data Facility cluster. This research was supported by a European Research Council (ERC) Starting Grant (No. 101041159).
Computational details
The starting point for constructing the position- and momentum-dependent exciton potential is a series of monolayer WS2 systems, each subjected to uniform strain by modifying the unit cell dimensions and internal atomic coordinates accordingly. Tensile strain is applied by enlarging (positive change) the unit cell, while compressive strain is applied by shrinking (negative change) the unit cell. These changes can be applied to a single in-plane crystal axis, resulting in uniaxial strain; to both in-plane crystal axes, resulting in uniform strain; or to the out-of-plane axis, resulting in out-of-plane strain. In this work we focus on the study of uniform strains as an illustrative example. The atomic geometry of the unit cells remains unchanged after their modification, since this procedure can induce coupling between different strain modes, which we aim to avoid in this study.
For each strained configuration, we perform Density Functional Theory (DFT) calculations, followed by a calculation of the quasiparticle (QP) corrections within many-body perturbation theory at the G0W0 (GW) approximation. The exciton band structure is then obtained by solving the Bethe-Salpeter equation (BSE), yielding the momentum-resolved exciton energies required for our dynamics model. Full computational details are given in the SI.

### v2

@@ Conclusions
To conclude, we presented a first-principles-based framework for modeling the ballistic time evolution of exciton distributions in inhomogeneously strained semiconductors, applied to strained monolayer WS2. Exploring idealized polynomial strain profiles, our results capture the qualitative impact of band curvature and energy gradients on exciton drift and diffusion patterns, in the absence of scattering and relaxation mechanisms. The model qualitatively reproduces the experimentally observed behaviors of drift toward strain maxima, anisotropic broadening, and negative-like diffusion. It also provides a physical basis for estimating the ballistic mean free time in idealized regimes. These results offer fundamental insight into how strain modulates exciton transport in 2D semiconductors and establish a foundation for ab initio-informed design principles in optoelectronic devices, where strain landscapes are engineered to control quasiparticle motion.
@@
Acknowledgments: We thank Moshe Harats, Lev Melnikovski, Diana Qiu, and Galit Cohen for valuable discussions. Computational resources were provided by the ChemFarm local cluster at the Weizmann Institute of Science and the Max Planck Computing and Data Facility cluster. This research was supported by a European Research Council (ERC) Starting Grant (No. 101041159).